OpenStudy (anonymous):

Okay, another hard one (for me at least). The function, p(d) = 1 + d/33, gives the pressure, in atmospheres (atm), at a depth d in the sea (d is in feet). Note that p(0)=1 atm, p(33) = 2, and so on. Find the pressure at 170 feet. Good luck with this one. I don't even understand the question! LOL

OpenStudy (anonymous):

For this one, d is the depth in feet, and p is the pressure. They are giving us a depth at 170 ft, and asking for the pressure. So we need to plug in d = 170 into the equation.

jimthompson5910 (jim_thompson5910):

p(170)=1+170/33 = 1+5.15151515151516 = 6.15151515151516

OpenStudy (anonymous):

\[p(170) = 1+\frac{170}{33} \] and jim already posted the answer.

OpenStudy (anonymous):

So the pressure at 170 feet is 1+ 170/33??

OpenStudy (anonymous):

You dont really have to understand what pressure is, or atm, or any of the technical jargon. Thats usually what confuses people. Just know you have a function, d is depth, p is pressure. They give me a depth, they want a pressure.

OpenStudy (anonymous):

Or is that what you have to simplify?

OpenStudy (anonymous):

well you would want to simplify it. Jim posted the answer.

jimthompson5910 (jim_thompson5910):

You have to simplify 1+170/33 to get the answer of 6.15151515151516, this is the pressure when you're 170 ft deep

OpenStudy (anonymous):

You could either enter it into a calculator or do it by hand. are you allowed to use calculators?

OpenStudy (anonymous):

Yes we are allowed but I don't have an algebraic calculator (yet).

jimthompson5910 (jim_thompson5910):

Use an online calculator then

jimthompson5910 (jim_thompson5910):

Google is a good one

OpenStudy (anonymous):

a normal calculator should be able to handle it. Just do: \[170\div33+1\]

OpenStudy (anonymous):

wouldn't it be 6.15?

OpenStudy (anonymous):

thats right :)

jimthompson5910 (jim_thompson5910):

type 1+170/33 into google to get your answer. http://www.google.com/search?q=1%2B170%2F33

OpenStudy (anonymous):

Whoo Hoo